# Impact of Data Distribution on Fairness Guarantees in Equitable Deep Learning

### v1

@@ 5 Conclusion
This work presents a theoretical analysis of the impact of data distributions on the fairness guarantees of deep learning models. Our analysis reveals that fairness is fundamentally constrained by the distributional differences between demographic groups, as demonstrated through our derived bounds on fairness errors, convergence rates, and group-specific risks. These theoretical insights are validated through extensive experiments on multiple datasets, where groups with feature distributions deviating significantly from the overall distribution consistently show lower performance, particularly across racial categories.

### v2

@@ 6 Conclusion
This work presents a theoretical analysis of how data distributional differences impact the fairness guarantees of deep learning models. Our results show that fairness is fundamentally constrained by inter-group shifts in feature means and covariances, as reflected in the derived bounds on fairness errors, convergence rates, and group-specific risks. Building on these insights, we proposed Fairness-Aware Regularization (FAR), a practical intervention that directly minimizes inter-group discrepancies in feature distributions. Extensive experiments across six large-scale datasets spanning medical images, natural images, tabular data, and natural language confirm our theory: demographic subgroups with larger distributional shifts consistently exhibit lower predictive performance, while incorporating FAR improves overall AUC, ES-AUC, and subgroup equity. Together, these findings bridge theoretical guarantees with algorithmic design, offering a principled and scalable pathway toward fairer deep learning models. Future research directions include extending the framework to multimodal foundation models and exploring adaptive strategies that tailor fairness interventions to specific subgroup characteristics.
Acknowledgments
This work was supported by NIH R00 EY028631, NIH R21 EY035298, NIH P30 EY003790, Research To Prevent Blindness International Research Collaborators Award, and Alcon Young Investigator Grant. We also acknowledge the generous funding resources provided by NYU Abu Dhabi with code AD131.